# Vision-Language Models can Identify Distracted Driver Behavior from Naturalistic Videos

### 04 Jan 2024 v3

@@ V-B Performance Comparison with Other Works
TABLE XIV: VideoCLIP at different sampling rates on SynDD1 dataset (with leaving a set of driver out)
@@
V-C Hyperparameters and Ablation Studies
1)  Sampling Rate
@@ V-C Hyperparameters and Ablation Studies
To see the effect of sampling rates on model performance, we test Multi-frameCLIP and VideoCLIP at various sampling rates on the SynDD1 dataset. Table XIII and Table XIV show the results. We maintain the same experimental settings and observe distinct trends across different sampling rates. The ViT-L/14-based Multi-frameCLIP achieves the peak Top-1 accuracy at a 1 FPS sampling rate with an accuracy of 70.36%. This was the highest accuracy achieved across all tested models and sampling rate settings for this experiment. In comparison, the AlexNet and VGG16 models got the lowest accuracy values. However, in Table XIV the VideoCLIP achieves the highest Top-1 accuracy at 20 FPS, with an accuracy of 81.85%. This contrast highlights the models’ sensitivity to sampling rate variations, suggesting that the optimal rate depends on the specific model and dataset characteristics. The changes from one frame to the next are more significant at a 1 FPS sampling rate, which helped the Multi-frameCLIP model make more accurate predictions. Therefore, 1 FPS is the optimal sampling rate for the Multi-frameCLIP model under the conditions of our study. On the other hand, as the average action duration for the SynDD1 dataset is $\sim$18.7 sec, a 20FPS rate results in a better temporal resolution, allowing the VideoCLIP model to capture more subtle changes and movements. If the sampling rate is too small, then the video clip will span more than the average action duration and confuse the model prediction. Therefore, the impact of the sampling rate would depend on the average action duration and the characteristics of the downstream tasks.
2)  Class-level Analysis
We illustrate the class-level F1-scores of Single-frameCLIP on the DMD and SAM-DD datasets in Table XII. In this table, we analyze the ten classes of the datasets to understand the most challenging actions to classify. The average F1-scores on DMD and SAM-DD datasets are 0.819 and 0.847, respectively. In addition, we demonstrate the confusion matrix of Single-frameCLIP on fold02 of the DMD dataset as an example in Fig. 8. We have found that the DMD classes such as “reaching to backseat”, “reaching to side” and “driving safely” exhibited lower F1-scores ($<$0.70), indicating that these categories pose greater challenges for the framework to classify correctly. These action classes appeared to generate similar image features, leading to a higher rate of misclassification. Additionally, we noticed that only two SAM-DD classes “head dropping” and “touching hair” got F1-scores below 0.70 among the ten classes. To further illustrate this, Fig. 9 provides visual examples of instances where the Single-frameCLIP model incorrectly classified actions, comparing the model’s outputs against the ground truth.
TABLE XV: Single-frame CLIP: Removing vs. merging classes results on SynDD1 dataset
Removing 5 3 0.64 0.71 0.84 0.82 0.80 - 0.763 $75.62\%$
Merging 6 3 0.51 0.50 0.82 0.75 0.88 0.73 0.698 $69.41\%$
Removing with
dashboard-view
5 1 0.77 0.87 0.89 0.94 0.89 - 0.873 $87.08\%$
TABLE XVI: Effect of Camera-view on SAM-DD dataset
@@
3)  Merging and Removing Classes
To fully understand the impact of the challenging classes on overall model performance, we pursue two additional experiments: merging and removing classes. Our initial focus was to explore how the overall model performance is influenced by either merging or removing classes that are more challenging to predict accurately. From our previous analysis, we noticed that the Single-frameCLIP model has difficulty in accurately predicting the “singing”, “eating” and “yawning” classes on the SynDD1 dataset compared to other classes. Table XV reflects our findings, which are categorized into three different experimental settings. The first row shows the outcome following the removal of the three challenging classes with three camera views. This modification resulted in an average F1 score of 0.763 and a mean accuracy of 75.62%. The second row illustrates the outcome of merging the three challenging classes with three camera views. This scenario led to a mean F1 score of 0.698 and an average accuracy of 69.41%. The third row provides the results obtained from an experiment in which the three challenging classes were removed and only the dashboard camera view was considered. This experimental setup led to an average F1 score of 0.873 and a mean accuracy of 87.08%. In conclusion, the optimization of the camera views, combined with the adjustment of the challenging classes, resulted in a significant improvement in the model’s performance.
4)  Effect of Camera Views
In this section, we investigate the effect of camera position on our framework’s performance. In Table XVI, we present the Top-1 accuracy for Single-frameCLIP and Multi-frameCLIP on ten distracted classes for the SAM-DD dataset, considering two distinct camera views. The outcomes suggest that the side view yields superior results when compared to the dashboard camera view. These results underscore the significance of the camera’s perspective in detecting driver distractions. The side camera view appears to provide the most effective angle for the SAM-DD dataset, capturing the driver’s movements and body maneuvers.
V-D Computational Complexity
TABLE XVII: VideoCLIP computation efficiency (Batch size=1)
@@ V-D Computational Complexity
In Table XVII, we present the details of the computational efficacy of the proposed VideoCLIP model. In terms of driver distraction recognition tasks, we compared the floating point operations (FLOPs) and the number of parameters (##\##Params.) of VideoCLIP with its counterparts. VideoCLIP demonstrates superior FLOPs efficiency and a relatively smaller parameter count ($<$10M). For example, in the SAM-DD dataset [4] setting, Swin-T gets 4.5G FLOPs and 28.3M trainable parameters. In [69] setting, VideoMAE operates at 57G FLOPs with 22M parameters, ResNet50 [42] uses 3.8 GFLOPs with 25.6M parameters for an input size of 224x224x3. Therefore, the computational performance of the VideoCLIP model coupled with a Top-1 accuracy of 81.85% makes it a resource-effective model in the driver distraction recognition task context. This efficiency is particularly valuable in real-time Intelligent Transportation applications like driver distraction recognition, where quick and accurate video processing is essential for safety and effectiveness. Moreover, the small parameter count also implies potentially lower memory requirements and faster training times, further enhancing its applicability in practical scenarios.
V-E Anecdotal Observation
@@
We demonstrate some miss-classified frames from the Single-frameCLIP model on DMD and SynDD1 dataset in Fig. 9 and the class-level confusion matrix of DMD classes in Fig 8. Analyzing the confusion matrix and the miss-classified frames, it is evident that the most challenging classes for the model occur where there is considerable feature overlap between classes that exhibit lower F1-scores. The middle column of Fig 9 shows the overlapping cues between two actions characterized by specific hand movements or facial expressions.
Figure 9: Some example frames that were misclassified by the Single-frameCLIP on DMD (top row) and SynDD1 (bottom row). These examples show evidence of how Single-frameCLIP fails to capture the temporal information only by looking at a single frame at a time.
@@
Fig. 10 presents the frame-by-frame analysis of video input for driverID 25470 from the SynDD1 dataset using the Multi-frameCLIP and VideoCLIP models. The depicted plot conveys the dynamic probabilities of predictions produced by the Multi-frameCLIP and VideoCLIP models over the entire span of action. The vertical axis shows the moving average of the computed probabilities in relation to the individual frame numbers, represented by the horizontal axis. We observe distinct fluctuations between high and low probability scores, aligning with frames where the model either successfully or unsuccessfully classified the given action. Our analysis focused on two distinct actions - adjusting hair and singing, at a sampling rate of 20 FPS. The representation of lower and higher probability scores from both models is clearly distinguished in this visualization. An important observation is the superiority of VideoCLIP’s prediction score over the entire course of action compared to that of the Multi-frameCLIP model. Additionally, we notice that VideoCLIP demonstrates a stable and consistent performance, capturing low-probability frames and maintaining a consistent score throughout the video. In contrast, the Multi-frameCLIP model appears to struggle at certain frames. Therefore, we conclude that VideoCLIP has the capability to efficiently capture temporal information from videos, significantly improving prediction accuracy while maintaining consistency. Moreover, the plot illustrates the superiority of both Multi-frameCLIP and VideoCLIP in extracting the temporal correlations among frames, thus providing insights into continuous actions that are challenging to get from isolated frames, as in the Single-frameCLIP scenario.
V-F Limitaions and Future directions
@@
Our algorithm shows promising results within its current setup, yet there is potential to further its adaptability and performance in varied real-world driving conditions across different unusual distraction categories. The variety of distractions that drivers encounter is vast and often unpredictable. Therefore, we need a sophisticated monitoring system to recognize the natural interactions between humans and vehicles and encourage safe driving practices. This system must be equipped to accurately identify the driver’s current state and assess the reliability of its predictions. Our current models were tested under controlled and available distracted settings from public datasets. The performance and efficacy in real-world applications, considering diverse hardware and environmental conditions, were not fully explored. To enhance the robustness and applicability of our algorithm, future work could involve training with additional data that cover a wider range of driving conditions - multiple distracted behaviors occurring simultaneously or alternately, unexpected or out-of-distribution behavior categories, occlusion events and so on. This would help in validating the algorithm’s effectiveness across various scenarios. Optimizing the algorithm for efficient performance across a range of hardware platforms, including low-power devices or with limited compute resources, could be a significant area of focus. This would make the system more viable for real-time applications in vehicles. Additionally, future research should also consider the ethical implications and privacy concerns associated with driver monitoring, ensuring compliance with relevant regulations and societal norms. Also, we plan to address the misclassification issues (discussed in V-E1) and enhance the reliability of the model predictions by implementing reliable uncertainty estimates [70]. This would be particularly beneficial in complex or ambiguous scenarios where driver behavior is not distinctly classifiable, thus allowing for a more nuanced understanding and interpretation of the model’s confidence measures. We believe that this addition will enhance the safety and reliability of the automated driver assistance systems by providing a more comprehensive and trustworthy analysis of driver behavior.
@@ VI Conclusion
In this paper, we have introduced a vision-language modeling framework aimed at categorizing distracted driving behaviors from naturalistic driving data. The performance of various CLIP-based frameworks has been evaluated on four common distracted driving datasets. Compared to the Zero-shotCLIP model, the Single-frameCLIP model benefited from task-based fine-tuning with significant improvement in accuracy while failing to capture the temporal dynamics of a video. Further experiments have shown that Multi-frameCLIP and VideoCLIP can leverage temporal information from video data, which allows for a better understanding of actions, events and sequences within a video. VideoCLIP and Multi-frameCLIP frameworks have outperformed Single-frameCLIP by a significant margin. Overall, our proposed frameworks exhibit state-of-the-art performance at distracted action classification and outperform traditional and most widely used models. This study has also evaluated the efficacy of different CLIP-model backbones, highlighting that a variety of factors, including variable training dataset size, sampling rates, and camera angles, can influence their performance. These observations can contribute to the development of more robust driver assistance systems to understand driver distractions. This study emphasizes the real-world applicability and generalizability of these models while also pointing out areas for improvement, such as handling unexpected and out-of-distribution driving behavior categories and fine-tuning hyperparameters. Several future works could extend and refine our current CLIP-based driver distraction recognition model. First, exploring the optimization of prompts for Vision-Language foundation models may enable the framework to extract more nuanced visual-textual relationships, leading to an improved classification. Second, the fusion of different camera views can be investigated to deal with occlusion, provide a more comprehensive picture of driver behavior and make accurate decision-making.
Moreover, the generalizability of our model across various scenarios and datasets presents another intriguing prospect. For instance, assessing whether the model could be fine-tuned on one dataset and subsequently deployed on another could further illustrate its versatility and effectiveness. In addition, we recognize the importance of empirical validation and testing our system under more challenging, unexpected, and out-of-distribution driving behavior categories. To this end, we are actively planning an extension of our current work. This next phase will include a series of experiments specifically designed to evaluate our framework’s response to real-world out-of-distribution actions. We also aim to extend the model’s testing to include a broader spectrum of actions, particularly focusing on those that are rare or challenging to classify. These tests are intended to validate our framework’s robustness and enhance its adaptive capabilities further. Finally, systematic evaluations under diverse environmental conditions, such as varying lighting conditions and data quality, could offer valuable insights into the model’s robustness and ability to function optimally under real-world driving circumstances.

### 21 Mar 2024 v4

@@ V-D Hyperparameters and Ablation Studies
TABLE XII: VideoCLIP at different sampling rates on SynDD1 dataset (with leaving a set of driver out)
@@
2)  Merging and Removing Classes
To fully understand the impact of the challenging classes on overall model performance, we pursue two additional experiments: merging and removing classes. Our initial focus was to explore how the overall model performance is influenced by either merging or removing classes that are more challenging to predict accurately. From our previous analysis, we noticed that the Single-frameCLIP model has difficulty in accurately predicting the “singing”, “eating” and “yawning” classes on the SynDD1 dataset compared to other classes. Table X reflects our findings, which are categorized into three different experimental settings. The first row shows the outcome following the removal of the three challenging classes with three camera views. This modification resulted in an average F1 score of 0.763 and a mean accuracy of 75.62%. The second row illustrates the outcome of merging the three challenging classes with three camera views. This scenario led to a mean F1 score of 0.698 and an average accuracy of 69.41%. The third row provides the results obtained from an experiment in which the three challenging classes were removed and only the dashboard camera view was considered. This experimental setup led to an average F1 score of 0.873 and a mean accuracy of 87.08%. In conclusion, the optimization of the camera views, combined with the adjustment of the challenging classes, resulted in a significant improvement in the model’s performance.
3)  Sampling Rate
@@
To see the effect of sampling rates on model performance, we test Multi-frameCLIP and VideoCLIP at various sampling rates on the SynDD1 dataset. Table XI and Table XII show the results. We maintain the same experimental settings and observe distinct trends across different sampling rates. The ViT-L/14-based Multi-frameCLIP achieves the peak Top-1 accuracy at a 1 FPS sampling rate with an accuracy of 70.36%. This was the highest accuracy achieved across all tested models and sampling rate settings for this experiment. In comparison, the AlexNet and VGG16 models got the lowest accuracy values. However, in Table XII the VideoCLIP achieves the highest Top-1 accuracy at 20 FPS, with an accuracy of 81.85%. This contrast highlights the models’ sensitivity to sampling rate variations, suggesting that the optimal rate depends on the specific model and dataset characteristics. The changes from one frame to the next are more significant at a 1 FPS sampling rate, which helped the Multi-frameCLIP model make more accurate predictions. Therefore, 1 FPS is the optimal sampling rate for the Multi-frameCLIP model under the conditions of our study. On the other hand, as the average action duration for the SynDD1 dataset is $\sim$18.7 sec, a 20 FPS rate results in a better temporal resolution, allowing the VideoCLIP model to capture more subtle changes and movements. If the sampling rate is too small, then the video clip will span more than the average action duration and confuse the model prediction. Therefore, the impact of the sampling rate would depend on the average action duration and the characteristics of the downstream tasks.
4)  Effect of Camera Views
TABLE XIII: Effect of Camera-view on SAM-DD dataset
@@
In this section, we investigate the effect of camera position on model performance. Table XIII presents the Top-1 accuracy for Single-frameCLIP and Multi-frameCLIP on the SAM-DD dataset, considering two distinct camera views. The outcomes suggest that the side-view yields superior results when compared to the dashboard camera view. These results underscore the significance of the camera’s perspective in detecting driver distractions. The side camera view appears to provide the most effective angle for the SAM-DD dataset, successfully capturing the driver’s movements and body maneuvers.
V-E Computational Complexity
TABLE XIV: VideoCLIP computation efficiency (Batch size=1)
@@ V-E Computational Complexity
In Table XIV, we present the details of the computational efficacy of the proposed VideoCLIP model. In terms of driver distraction recognition tasks, we compared the floating point operations (FLOPs) and the number of parameters (##\##Params.) of VideoCLIP with its counterparts. VideoCLIP demonstrates superior FLOPs efficiency and a relatively smaller parameter count ($<$10M). For example, in the SAM-DD dataset [18] setting, Swin-T gets 4.5G FLOPs and 28.3M trainable parameters. In [79] setting, VideoMAE operates at 57G FLOPs with 22M parameters, ResNet50 [49] uses 3.8 GFLOPs with 25.6M parameters for an input size of 224x224x3. Therefore, the computational performance of the VideoCLIP model coupled with a Top-1 accuracy of 81.85% makes it a resource-effective model in the driver distraction recognition task context. This efficiency is particularly valuable in real-time Intelligent Transportation applications like driver distraction recognition, where quick and accurate video processing is essential for safety and effectiveness. Moreover, the small parameter count also implies potentially lower memory requirements and faster training times, further enhancing its applicability in practical scenarios.
V-F Anecdotal Observation
@@
We demonstrate some miss-classified frames from the Single-frameCLIP model on the DMD and SynDD1 dataset in Fig. 6. Analyzing the confusion matrix and the miss-classified frames, it is evident that the most challenging classes for the model occur where there is considerable feature overlap between classes that exhibit lower F1-scores. The middle column of Fig 6 shows the overlapping cues between two actions characterized by specific hand movements or facial expressions.
Figure 6: Some example frames that were misclassified by the Single-frameCLIP on DMD (top row) and SynDD1 (bottom row). These examples demonstrate how Single-frameCLIP fails to capture the temporal information only by looking at a single frame at a time.
@@
Fig. 7 presents the frame-by-frame analysis of driverID 25470 from the SynDD1 dataset using the Multi-frameCLIP and VideoCLIP. It demonstrates the dynamic probabilities of predictions produced by the Multi-frameCLIP and VideoCLIP models over the entire span of an action. The vertical axis shows the moving average of the computed probabilities against the individual frame numbers, represented by the horizontal axis. We observe distinct fluctuations between high and low probability scores, aligning with frames where the model either successfully or unsuccessfully classified the given action. Our analysis focuses on two distinct actions - adjusting hair and singing, at a sampling rate of 20 FPS. The representation of lower and higher probability scores from both models is clearly depicted in this plot. An important observation is that the VideoCLIP demonstrates a stable and consistent prediction score throughout the entire video compared to that of the Multi-frameCLIP. Therefore, we conclude that VideoCLIP has the capability to efficiently capture temporal information from videos, significantly improving prediction accuracy while maintaining consistency. Moreover, the plot illustrates the superiority of both Multi-frameCLIP and VideoCLIP in extracting the temporal correlations among frames, thus providing insights into continuous actions that are challenging to get from isolated frames, as in the Single-frameCLIP scenario.
V-G Limitaions and Future directions
@@
Our algorithm shows promising results within its current setup, yet there is potential to further its adaptability and performance in varied real-world driving conditions across different unusual distraction categories. The variety of distractions that drivers encounter is vast and often unpredictable. Therefore, we need a sophisticated monitoring system to recognize the natural interactions between humans and vehicles and encourage safe driving practices. This system must be equipped to accurately identify the driver’s current state and assess the reliability of its predictions. Our current models were tested under controlled and available distracted settings from public datasets. The performance and efficacy in real-world applications, considering diverse hardware and environmental conditions, were not fully explored. To enhance the robustness and applicability of our algorithm, future work could involve training with additional data that cover a wider range of driving conditions - multiple distracted behaviors occurring simultaneously or alternately, unexpected or out-of-distribution behavior categories, occlusion events and so on. This would help in validating the algorithm’s effectiveness across various scenarios. Optimizing the algorithm for efficient performance across a range of hardware platforms, including low-power devices or with limited compute resources, could be a significant area of focus. This would make the system more viable for real-time applications in vehicles. Additionally, future research should also consider the ethical implications and privacy concerns associated with driver monitoring, ensuring compliance with relevant regulations and societal norms. Also, we plan to address the misclassification issues and enhance the reliability of the model predictions by implementing reliable uncertainty estimates [80]. This would be particularly beneficial in complex or ambiguous scenarios where driver behavior is not distinctly classifiable, thus allowing for a more nuanced understanding and interpretation of the model’s confidence measures. We believe that this addition will enhance the safety and reliability of the automated driver assistance systems by providing a more comprehensive and trustworthy analysis of driver behavior.
@@ VI Conclusion
In this paper, we introduce a vision-language framework for recognizing distracted driving behaviors from naturalistic driving videos. Our findings reveal that task-based fine-tuning enhances the Single-frameCLIP model’s accuracy, while Multi-frameCLIP and VideoCLIP models can leverage the temporal contexts of a video and significantly outperform the frame-based models. Our proposed frameworks exhibit state-of-the-art performance at distracted action recognition and outperform the recent benchmark models. This study emphasizes the real-world applicability and generalizability of these models while also pointing out areas for improvement. Future work will focus on optimizing visual-textual relationships, integrating multiple camera views for comprehensive behavior analysis, and testing the model’s generalizability across various real-world scenarios. Moreover, the generalizability of our model across various scenarios and datasets presents an intriguing prospect. In addition, we recognize the importance of empirical validation and testing our system under more challenging, unexpected, and out-of-distribution driving behavior categories. To this end, we are actively planning an extension of our current work and aiming to extend the model’s testing to include a broader spectrum of actions, particularly focusing on those that are rare or challenging. These tests are intended to validate our framework’s robustness and enhance its adaptive capabilities further. Finally, systematic evaluations under diverse environmental conditions, such as varying lighting conditions and data quality, could offer valuable insights into the model’s robustness and ability to function optimally under real-world driving circumstances.